OpenStudy (anonymous):

Choose the missing number to make equivalent fractions. 1/4 = blank/16 A. 2 B. 4 C. 6 D. 8

OpenStudy (anonymous):

16/4=4 so B.

OpenStudy (anonymous):

B because 1/4 is .25 and 4/16 is also .25

OpenStudy (charlotte123):

the answer would be 4 :D because the denominator was multiplied by 4 to get 16, so multiplying the numerator by 4, u would get 4 ^_^ 1/4 = 4/16 :D

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):